Instrument Log

25MT ECOR C1                 CF Lamont, OK

Date

Operator

Comments

4/20/01

ZW

PM and disk change out. The 60M ECOR sonic would not respond to commands from the laptop computer. Power cycled the sonic at 1430 GMT. Communications were restored. The 60M Hygrometer also is working fine, however there are no instruments on the 25 M boom at this time. RTS 1500 GMT.

3/22/01

MK

No PM performed. Instruments not installed at 25M level.

2/1/2001

MR/KT

Lowered the SW (B) side of the 60M tower from 2140 GMT to 2125 GMT to secure a loose hygrometer cable at the 60M level. All instruments on the SW side of the tower were affected during this time.

6/23/00

ZW

The 60M ECOR was missing it's DCU. The C1 ECOR *DCU was picked up at storage, & installed at 2000 GMT. The two disks for the 25M level, & 60M level both stopped spinning. The new DCU communicated with the laptop however there are only instruments on the 60M level. The 60M sonic, & hygrometer both had good readings. There are no spare C1 sonics in spares at this time. When a spare C1 sonic becomes available it, and a hygrometer will be installed at the 25M level. RTS 2100 GMT. *New - 940101

8/13/99

ZW

The 25M SONIC, and 25M Hygrometer are missing at this time.

7/30/99

MK

The 25M SONIC, and 25M Hygrometer are missing at this time.

7/16/99

MK

The 25 M Hygrometer would not respond to operator commands. Lowered tower at 1430 GMT and R/R Hygrometer, but still no response. R/R Hygrometer with another spare, still no response. Checked the continuity of the digital data cable from the DCU to the Hygrometer, no problems found. Checked DC power going in to the Hygrometer, only measured 8.45 volts DC inside Hygrometer . Unlatched DC cable from Hygrometer box and measured 12 volts DC coming from the SONIC. Used external DC power supply to power up Hygrometer so we could get the full 12 volts DC, all communications with the Hygrometer were restored, tried this with the other two Hygrometers, all worked normal. The 110 volt AC to 12 volt DC power supply inside the SONIC is not strong enough to properly power up the Hygrometer. Removed the *SONIC and **Hygrometer from the 25 M until another Sonic becomes available. *Old - 940404. **Old - 940216

7/2/99

ZW

The C1 60/25 meter ECOR has been down because there was no C1 DCU in spare. The C1 ECOR *DCU came in so Site Ops took it up to the tower, and installed it at 1720. Site Ops also installed a new **SONIC and a ***Hygrometer on the 25M platform. The 60M and 25M meter SONIC's worked fine, however the 25M and 60M Hygrometer would not communicate with the laptop computer. Site Ops power cycled the system several times with no success. RTS 2020 GMT. *New - 960201 **New - 940404 ***New - 940216

5/19/99

5/20/99

CAM/DH

Dick Hart and Gary Zimmerman of ATI on site to T/S persistent 25/60M ECOR problems. Lowered the tower booms at 1755 GMT on 5/19/99. There were currently no sensors installed at the 25M level but Gary determined from the 60M sensors that the problems we have been experiencing are most likely due to the DCU software. Removed the *DCU and will return to ATI for diagnostics and software corrections. Attempted to install the spare DCU but encountered similar problems. Spare DCU will also be returned to ATI. Raised booms at 1755 GMT on 5/20/99 with no 25M ECOR installed. *Old - 940101.

4/30/99

MK

A Hygrometer and Sonic need to be installed at the 25M level. Lowered tower at 1515 GMT. Installed Sonic sensor and Hygrometer sensor. Powered both sensors on using extension cables. The Hygrometer would work , but the Sonic would not respond to operator commands. Power cycled several times, but still no response. The error message 1234 PUDIAG was the only response. Removed the Sonic and Hygrometer sensors until another spare Sonic sensor is available. The bad Sonic sensor S/N is 940404.

2/12/99

MK

The 25M SONIC is not responding to operator commands. Power cycled several times at 1635 GMT, still no response, lowered tower at 1645 GMT and removed *25M SONIC so it can be sent off for repair. No spares at this time. The 25M hygrometer is not installed at this time. *Old - 940404

1/29/99

ZW

PM and disk change out. The 25M SONIC is not operational and the 25M Hygrometer is missing. Due to lack of spare sensor no corrective action taken at this time. 60M SONIC, Hygrometer are working in fine order. Status ok. RTS 1515 GMT.

1/15/99

ZW

PM and disk change out. The 60M SONIC and 60M Hygrometer are working fine. The 25M SONIC, and 25M Hygrometer are down at this time and will be down until spares become available. RTS 1445 GMT.

12/30/98

ZW

PM and disk change out. The 25M SONIC, and 60M Hygrometer would not respond to operator commands from the laptop computer. The 25M Hygrometer is missing at this time. Power cycled the 25M SONIC, and the 60M Hygrometer several times however this did not fix this problem. Due to lack spare sensors no more corrective action was taken at this time. The 60M SONIC is the only sensor working at this time. RTS 2000 GMT.

12/17/98

ZW

PM & disk change out. The C1 ECOR 1a SONIC, and 1c Hygrometer would not respond to operator commands. Power cycled the 1a SONIC, and 1c Hygrometer at 1940 GMT. This did not fix the problem. Because of no spare C1 SONIC there was no more corrective action taken at this time. RTS 1955 GMT

12/7/98

ZW

PM and disk change out. The 1a SONIC would not respond to operator command’s. The 2b SONIC is working fine. The 1c Hygrometer would not respond to operator command’s. The 2d Hygrometer is also working fine at this time. Because of lack of spare C1 SONIC or Hygrometer no corrective action was taken at this time.

10/22/98

ZW

PM. The 25M ECOR is down at this time because of no spare parts.

10/16/98

SS/MK

Work order requested that the hygrometer be installed and tested. Lowered the "B" side of the tower and installed a spare hygrometer and tested. Found that the DCU would not communicate with the hygrometer. Installed a second spare hygrometer and still could not communicate with it. Tested one of the spare hygrometers with the E14 ECOR and found that the hygrometer is good. Removed the hygrometer from the 25 meter tower and raised the instruments back up. Hygrometers were returned to spares. RTS at 1630 GMT.

10/15/98

MW/MR

There was no sonic sensor connected at the 25M level. Attempted to get the 25M ECOR instrument suite fully operational this morning by installing a new *sonic sensor with the existing hygrometer. Lowered the boom at 1330 GMT. Installed sonic sensor on the 25M level. Was able to communicate with the sonic sensor but could not establish communications with the existing hygrometer. Removed **hygrometer from service. Attempted to replace the hygrometer but the spare hygrometer electronics box sensor connector was missing its locking lugs so the sensor connector could not be properly secured to the electronics box. Will place another hygrometer into service when spares become available. Raised the boom at 1505 GMT. Returned to service with sonic sensor only. *New - 940404 **Old - 960202

10/14/98

MW/MR

The instrument mentor wanted to know the status of this instrument. The C1 DCU is fully operational. The 25M ECOR sensor suite contains no sonic sensor. There is a hygrometer on the 25M boom but it is non-operational since it derives its power from the sonic sensor. Checked spare storage and located another C1 sonic sensor. Did not lower the booms today due to high winds. Will lower the booms first thing tomorrow morning and install the spare sonic sensor on the 25M level. Replaced the 25M Bernoulli disk at 1955 GMT since there was only 9.6MB of free disk space left.

9/25/98

ZW

PM. The 25M ECOR is down at this time.

8/27/98

ZW

PM was not done at this time because the 25M ECOR is down at this time.

7/17/98

MR/KT

The 25M Sonic has not been operational since the 6/18 PM but had not yet been replaced due to lack of spares. Removed the *25M Sonic while the booms were lowered for T/H calibrations at 1445 GMT on 7/17/98. Will replace when spare becomes available. *Old - 940404

7/16/98

7/17/98

CAM/DC

All tower booms were lowered from 1458 GMT on 7/16/98 until 1513 GMT on 7/17/98 for 25M and 60M T/H calibrations. The booms were powered at ground level during that time.

7/1/98

ZW

PM and disk change out. The 1a SONIC and the 2d Hygrometer would not respond to operator commands. Power cycled the 1a SONIC and the 2d Hygrometer. the SONIC and Hygrometer would not respond to operator commands. Power cycled the ECOR system at 1945 GMT. The a1 SONIC and 2d Hygrometer still would not respond to operator commands form the laptop computer. Because of no spare sensors no more corrective action was taken. RTS 2000 GMT.

6/23/98

MW

SDS operations reported a loss of communications with all of the 60M tower instruments beginning at 1004 GMT today. Power cycled the tower COS at 1404 GMT. Communications restored after the power cycle. RTS at 1405 GMT.

6/18/98

ZW

PM and disk change out. The 1a SONIC would not respond to operator commands from the laptop computer, however the 1c Hygrometer was working fine. Power cycled several times with no success. Because of lack of spare C1 SONIC no further corrective action was taken at this time. RTS 1945 GMT.

6/4/98

ZW

PM and disk change out. The 25M SONIC is working in good order, however the 25M Hygrometer would not respond to operator commands from the laptop computer. Power cycled the Hygrometer at 2000 GMT. Still could not communicate with the Hygrometer. Due to lack of spares no more corrective action taken at this time. RTS 2015 GMT.

5/21/98

ZW

PM and disk change out. The DCU would not respond to operator commands from the Laptop computer. Power cycled the DCU at 1845 GMT. The communications with the DCU were restored. Status ok. RTS 1900 GMT.

5/7/98

ZW

PM and disk change. SONIC, and Hygrometer are working fine. Status ok. RTS 1925 GMT.

4/23/98

ZW

PM and disk change out. The 25M SONIC, and Hygrometer are running fine. RTS 1842 GMT.

4/10/98

ZW

PM and disk change out. Status ok. The 25M SONIC, and Hygrometer both are operating fine. RTS 1340 GMT.

3/26/98

ZW

PM and disk change out. The 25M SONIC is the only sensor operational at this time. Due to lack of spare 60M SONIC which should be back any time from Applied Technologies.

3/12/98

ZW

The 25M Hygrometer was not installed because the C1 SONIC was removed for repairs. The C1 SONIC is required because the power supply for the hygrometer is in the C1 SONIC. The C1 SONIC was returned from Applied Technologies after repairs. Site Ops installed *Sonic and **Hygrometer. Status ok. RTS 1730 GMT. *New - 940404. **New - 960202.

2/27/98

ZW

PM. The 25M Hygrometer, and the SONIC are still missing at this time.

2/13/98

ZW

The PM on the C1 ECOR was yesterday. It was discovered at this time that the DCU would not respond to operator commands. Attempts to power cycle the DCU would not work because the Bernoulli disk would not stop spinning. The 9 pin connector where the laptop connects to the DCU also needs to be replaced. At 1930 Site Ops powered off the C1 ECOR DCU. R/R *DCU with one just back from Applied Technologies INC. The new Hygrometer worked fine. RTS 2030 GMT * Old - 960201 * New - 940101

2/12/98

ZW

PM. No disk change because the 25M Hygrometer, and SONIC are missing at this time.

1/29/98

ZW

PM. The 25M SONIC, and Hygrometer are missing at this time. Removed disk but did not replace it at this time.

1/27/98

MW

The 'U' signal of the 25M sonic sensor was showing a steady 00.06 and was not changing. In comparison, the 60M sonic sensor was fluctuating around 6.XX or 7.XX for the same signal. Lowered the tower booms at 1755 GMT. Removed the 25M *sonic sensor from service. Also removed the 25M **hygrometer sensor from service since it gets its power from the sonic sensor. Do not have a spare C1 sonic sensor to place in service at this time. Raised booms at 1840 GMT. RTS. * Old – 940404 ** Old – 960203

1/16/98

ZW

PM and disks were changed out. RTS 1450 GMT.

12/31/97

ZW

PM and Disk change. Power cycled the ECOR at 1900 GMT. Status ok. RTS 1920 GMT.

12/12/97

ZW

PM. The C1 ECOR *DCU Bernoulli disk drives would not stop spinning. This is required to communicate with the DCU. R/R the disk drives which did not work. Removed the CPU board from the DCU which was operating the C1 ECOR at the time S/N 940101, and placed the CPU board in another double drive DCU S/N 960201 which came from the ECOR at E24 Cyril . The DCU From E24 Cyril S/N 960201 is working in the C1 ECOR at this time. *Old - 940101 *New - 960201

12/8/97

ZW

The SM122 Power Distribution Board was removed from the 25M ECOR SONIC by the Mentor, and sent back to Applied Technologies for repair. The SM122 Power Distribution Board was returned from Applied Technologies. Site Ops installed the SM122 Board on the 25M SONIC at 1900 GMT. RTS 2000 GMT.

11/5/97

CAM

After installing a new DCU on 10/24/97, the 25/60m ECOR could not be monitored. The Bernoulli disks spun continuously. Due to no spare DCU's and mentor being expected on site 11/4, the system was left as is for mentor to examine. Mentor arrived on 11/4 and the booms were lowered for maintenance. As expected, the system could not be monitored. Installed a video card in the DCU and hooked it up to a monitor. By monitoring the boot up sequence and diagnostics, it was determined that the DCU failed while trying to read drive D which is one of the Bernoulli drives. No spare Bernoulli drives on hand at this time. System remains off line until a spare is obtained.

11/4/97

CAM

The new 25m Sonic electronics that were installed on 10/24/97 were suspected to be faulty. The 25m GFCI would not stay reset since the electronics were installed. Instrument mentor was expected on site 11/4 so the power to the 25m level was shut off and the electronics left in place for mentor to examine. Mentor arrived on 11/4 and the booms were lowered for maintenance. T/S found that the power supply for the hygrometer electronics, which is an integral part of the sonic electronics, had failed. There was a dead short between the AC common and AC ground inputs. Removed the power supply to return to vendor for repair or replacement.

10/24/97

ZW

PM. The C 1 ECOR was down because the DCU was sent back to Applied Technologies to be fixed. Received the *DCU back and reinstalled. However the problem remains the same. The DCU will not communicate with the laptop computer. The Bernoulli disk for the 25 meter, and the Bernoulli disk for the 60 meter will not stop spinning. Communication with the DCU can not be established until the disks stop spinning. No corrective action taken at this time site ops will bring back the only other DCU with two Bernoulli disk drives for a possible alternate for the C 1 ECOR. The 25 meter ECOR was missing its SONIC. Replaced missing **SONIC. The 25 meter ECOR was missing its Hygrometer. Replaced missing ***Hygrometer. RTS 1950 GMT. *New - 940101 ** New - 940404 *** New – 960203

10/10/97

ZW

This instrument is not operational at this time.

9/12/97

ZW

PM and removed disks. The DCU would not respond to operator commands from the laptop computer. Power cycled the DCU several times to no avail. Removed *DCU to send in for repair. This instrument is down until a spare is available. * Old - 940101

8/15/97

ZW

PM and disks change out. The 25 meter SONIC and Hygrometer are missing. No corrective action taken due to lack of spare. The 60 meter SONIC and Hygrometer are working fine. RTS 1350 GMT.

8/14/97

GRS

After calibrations and repairs to the tower, the 25 boom was raised to its proper height at 1602 GMT. NOTE: There is no ECOR instruments at the 25m height.

8/13/97

GRS

Craig Webb installed the new plate on the lower part of the trolley. Calibrations of the temp/hum sensors caused the tower to remain down until the calibrations were complete.

8/5/97

MW

Was not able to communicate with the 25M Sonic sensor using a laptop computer. Powered the DCU off at 1758 GMT. Lowered the 25M boom at 1803 GMT. Removed 25M *Sonic and 25M **Hygrometer sensors from service. Did not raise the 25M boom due to a stress fracture in the aluminum plate where the bottom u-bolt connects the boom to the elevator cable. * Old - 940404 ** Old - 960202

7/18/97

ZW

PM and disks were changed. Power cycled DCU at 1730 GMT. Communications restored. The 1A SONIC is not working at this time. No spare C1 SONIC at this time. Other than that Status is ok. RTS 1740 GMT.

7/3/97

ZW

PM and disks change. SONIC 1A is bad no spare at this time. Other than that status ok. RTS 1530 GMT.

6/20/97

ZW

PM and disk change. SONIC A1 would not respond to operator commands. Power cycled system at 1600 GMT with no positive results. Status other than that ok. RTS at 1615 GMT.

6/5/97

ZW

PM and disk change. Power cycled DCU at 1910 GMT. Power cycled SONIC 1A, 2B, and Hygrometer 1C, 2D at 1920. SONIC 1A is still bad however there are no spare at this time. Other than that status ok. RTS 1940 GMT.

5/23/97

ZW

PM and disks changed. SONIC A1 is bad however there are no spare at this time. Other than that Status ok. RTS 1330 GMT.

5/9/97

ZW/MK

PM and disk change out. Power cycled DCU at 1437 GMT. SONIC A (25M) was giving bad readings power cycled SONIC A at 1451 GMT to no avail. No spare at this time. Other than SONIC A Status is ok. RTS 1506 GMT.

5/01/97

CM/MW

Lowered boom at 1720 GMT. *Installed repaired DCU and performed PM. Set DCU clock at 1835 GMT. Time delta was -00:04:17. Hygrometer Q reading high resulting in RH of near 80 percent when actual humidity was near 50 percent. No spare hygrometer at this time. Raised boom and RTS at 1920 GMT. * New - 940101

2/18/97

MDW/CAM

Further T/S suggested DCU was bad. *Removed DCU and returned to vendor for repair. * Old - 940101

2/11/97

CM/MS

T/S found that the 60M Sonic power supply was bad, thus tripping the GFCI. 25M Sonic and Hygrometer tested O.K. After changing the Bernoulli disk, communications could not be reestablished and there was no Bernoulli disk activity. Raised boom at 2120 GMT. T/S to continue.

2/10/97

MW/ZW

Lowered booms at 2000 GMT to install repaired * Sonic and ** Hygrometer. After sensors were installed the "B Upper" GFCI would not stay reset. Removed power and will T/S tomorrow. * New - 940404 ** New - 960202

10/21/96

MDW

Removed *hygrometer and **sonic. * Old - 940211 ** Old - 940404

9/30/96

MDW

PM and disk change. Could not communicate with the Hygrometer and Sonic. Power cycled the entire system at 19:46 GMT. Changed Bernoulli from 9/13/96 to 9/30/96. No communications at 20:00 GMT. Power cycled at 20:00 GMT. RTS at 20:15 GMT. Status O.K.

9/10/96

DKB/CAM

1942 GMT Lowered boom for 60 T/H PM. RTS at 2300 GMT.

8/12/96

CAM/MDW

1815 GMT. PM and data collection. Could not communicate with 25M sonic. Power cycled DCU and sensor by lowering boom slightly and then raising again. RTS status O.K. at 1830 GMT.

8/09/96

CAM/KDT

2010 GMT. T/S of 25M sonic problem. Lowered boom to re-seat connections and power cycled DCU. Still unable to communicate with sonic. Had to abandon T/S due to approaching t-storm. T/S to continue.

8/08/96

CAM/GRS

Lowered 25M ECOR at 2023 GMT. *Replaced sonic and tested O.K. on ground. Raised boom at 2050 GMT. Monitored again and sonic now read all -99’s. RTS at 2115 GMT as is. T/S to continue. * new - 940404

7/23/96

CM/GS

25 hygrometer power supply inside sonic box failed. * Removed sonic sensor and sent to vendor. 25M ECOR inop until sonic is repaired. * old - 940404

7/22/96

CM/MW

2000 GMT Lowered tower at 2000 GMT. T/S of hygrometer indicates blown fuse. R/R fuse but electronics still inop. * R/R hygrometer New hygrometer would not work also. T/S indicates no DC power from sonic box to hygrometer box. Raised tower and RTS at 2050 GMT with inop hygrometer. * old - 940209 * new - 940211

7/19/96

MW

PM. R/R Bernoulli disk. Could not communicate w hygrometer. Power cycled hygrometer but did not fix problem. RTS at 2130 GMT w/ bad hygrometer.

7/8/96

CAM

PM & Data collection.

7/3/96

CAM/MW

Special visit. Lowered tower at 1925 GMT to T/S Hygrometer. * R/R Hygrometer and tested OK at 2020 GMT. * old - 940214 * new - 940209

6/24/96

MW/ZW

PM & Data collection. Added water to the cool cell. Noticed that the Hygrometer H and Q values are reading out of range. Will T/S at a later date. RTS at 1815 GMT.

6/15/96

MW/ZW

*R/R hygrometer. * Old - 940217 * New - 940214

6/7/96

MW/ZW

PM & Data collection. RTS at 1800 GMT.